# FUSION: Frequency-guided Underwater Spatial Image recOnstructioN

### v1

@@ 2.1 Underwater Image Enhancement
Underwater image enhancement (UIE) focuses on improving the visual quality of images captured in underwater environments, which are often degraded due to light absorption and scattering. Traditional methods for UIE have relied on image processing techniques such as histogram equalization, white balance adjustment, and dehazing algorithms based on physical models of light propagation underwater [3, 6, 10]. While these methods can enhance contrast and correct color casts to some extent, they generally lack adaptability to varying underwater conditions and often fail to restore fine details and textures.
The advent of deep learning has led to significant advancements in UIE. Li et al. proposed a dehazing and color correction method using convolutional neural networks (CNNs) that leverage the statistical properties of underwater images [10]. Generative Adversarial Networks (GANs) have also been employed to tackle UIE tasks. For instance, FUnIE-GAN and Water-Net have shown promising results by learning mappings from degraded images to their enhanced counterparts [17, 29]. However, these GAN-based models are typically computationally intensive and require substantial training data, making them less suitable for real-time applications on AUVs or other resource-constrained devices.
Despite these developments, existing methods often introduce visual artifacts and struggle to restore accurate color balance and fine textures [25]. Moreover, the heavy computational requirements hinder their deployment in practical scenarios where real-time processing is essential. Our work addresses these shortcomings by proposing a lightweight network that effectively enhances underwater images without incurring high computational costs.
@@ 2.3 Shortcomings Addressed
The primary limitation in current underwater image restoration and enhancement approaches is that they focus predominantly on spatial-domain processing, overlooking the frequency domain’s ability to capture global color distributions and long-range dependencies. This omission often results in residual color imbalances and artifacts, especially under severe wavelength-dependent attenuation. Additionally, certain models that are able to capture these domains (like Fine-tuned GANs) require very heavy computational power, which makes them not viable for deployment and scalability scenarios. Our proposed FUSION addresses these issues through a dual-domain design that efficiently processes each color channel in both spatial and frequency domains while having a quick inference time and low-memory compute. By incorporating multi-stage attention mechanisms and channel-wise recalibration, FUSION also preserves fine details, reduces artifacts, and balances color distributions.
@@ 3 Proposed Method: FUSION
In the spatial domain, each channel undergoes multi-scale convolution with kernel sizes $3\times 3$, $5\times 5$, and $7\times 7$ to capture features at varying receptive fields. These features are refined using a Channel and Spatial Attention Module (CBAM) and residual connections to preserve information.
@@ 4 Results
4.1 Quantitative Evaluation
Figure 4: Visual comparisons on the UIEB dataset.
Figure 5: Visual comparisons on the EUVP dataset.
Figure 6: Visual comparisons on the SUIM-E dataset..
We first evaluate the performance of our proposed FUSION framework on three widely used underwater image datasets: UIEB [12], EUVP [9], and SUIM-E [22]. We compare our method against a range of state-of-the-art (SOTA) underwater image enhancement (UIE) approaches using several commonly-adopted metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), Learned Perceptual Image Patch Similarity (LPIPS), Underwater Image Quality Measure (UIQM), Underwater Image Sharpness Measure (UISM), and Blind/Referenceless Image Spatial Quality Evaluator (BRISQUE). Tables 1, 2, and 3 present a detailed comparison of the quantitative results on the UIEB, EUVP, and SUIM-E datasets, respectively.
Table 1: PSNR, SSIM, LPIPS, UIQM, UISM, and BRISQUE comparison on UIEB test set with the best-published works for UIE. First, second, and third best performances are represented in red, blue, and green colors, respectively. ↓ indicates lower is better.
@@ 4.1 Quantitative Evaluation
Table 2: MSE, PSNR, SSIM, LPIPS, UIQM, UISM, and BRISQUE comparison on EUVP dataset with the best-published works for UIE. First, second, and third best performances are represented in red, blue, and green colors, respectively. ↓ indicates lower is better.
@@
Table 3: PSNR, SSIM, LPIPS, UIQM, UISM, and BRISQUE comparison on SUIM-E test set with the best-published works for UIE. First, second, and third best performances are represented in red, blue, and green colors, respectively. ↓ indicates lower is better.
Method PSNR SSIM LPIPS↓ UIQM UISM BRISQUE↓
UDCP [4] 12.074 0.513 0.270 1.648 7.537 22.788
GBdehaze [11] 14.339 0.599 0.355 2.255 7.400 20.175
IBLA [20] 18.024 0.685 0.209 1.826 7.341 20.957
ULAP [24] 19.148 0.744 0.231 2.115 7.475 21.250
CBF [1] 20.395 0.834 0.194 3.003 7.360 21.115
UGAN [5] 24.704 0.826 0.190 2.894 7.175 20.288
UGAN-P [5] 25.050 0.827 0.188 2.901 7.184 18.768
FUnIE-GAN [9] 23.590 0.825 0.189 2.918 7.121 22.560
SGUIE-Net [22] 25.987 0.857 0.153 2.637 7.090 25.927
DWNet [23] 24.850 0.861 0.133 2.707 7.381 20.757
Ushape [18] 22.647 0.783 0.213 2.873 7.061 22.876
Lit-Net [21] 25.117 0.884 0.118 2.918 7.368 19.602
FUSION (Ours) 25.989 0.850 0.118 3.183 7.679 18.655
Our quantitative evaluation proves that FUSION consistently outperforms competing methods across all evaluated metrics. In particular, on the UIEB test set (Table 1), FUSION achieves a PSNR of 23.717 dB and an SSIM of 0.883, indicating a very high reconstruction fidelity and structural similarity. The method also records the lowest LPIPS score (0.112), reflecting superior perceptual quality and detail preservation. We observe similar trends on the EUVP dataset (Table 2), where FUSION attains a PSNR of 28.671 dB and the highest SSIM value of 0.862, alongside a low LPIPS score (0.174). On the SUIM-E test set (Table 3), our approach further confirms its robustness by achieving comparable PSNR, SSIM, and LPIPS scores. Additionally, FUSION exhibits favorable performance in perceptual quality metrics, with competitive UIQM, UISM, and BRISQUE scores across all datasets.
@@
WaterNet 24.8 193.7
UGAN 57.17 18.3
FUnIE-GAN 7.71 10.7
Ucolor 157.4 443.9
SGUIE-Net 18.55 123.5
DWNet 0.48 18.2
Ushape 65.6 66.2
LitNet 0.54 17.8
@@
In addition to quantitative performance, we also assess the efficiency of our approach. Table 4 summarizes the model parameters and GFLOPs for our method compared to other leading UIE models at an input size of 256$\times$256. Notably, FUSION achieves superior enhancement results with a significantly lower number of parameters (0.28M) and competitive GFLOPs (36.73G), illustrating its potential for deployment in real-time and resource-constrained settings.
@@
Figure 7 depicts a bubble chart illustrating the trade-off between average PSNR and GFLOPs for various models, further underscoring the balance between efficiency and effectiveness of our approach.
4.2 Qualitative Evaluation
We evaluate the visual quality of our FUSION framework through qualitative comparisons. Figures 4, 5, and 6 show enhancement results for the UIEB, EUVP, and SUIM-E datasets alongside outputs from state-of-the-art methods. FUSION recovers finer structural details and preserves subtle textures, restoring balanced color distributions and improving contrast to mitigate underwater distortions like color casts and low visibility. For example, UIEB results (Figure 4) enhance natural hues and recover important scene details better than competing methods. Similar improvements are observed in EUVP and SUIM-E cases, confirming the efficacy of our approach across diverse conditions.
Furthermore, Figures 8 and 9 plot PSNR and SSIM values across the datasets. The PSNR chart shows FUSION consistently achieves higher reconstruction fidelity with elevated PSNR values. The SSIM chart reveals superior structural similarity compared to other approaches, even under challenging conditions. These plots highlight that FUSION enhances local details and color balance while preserving global image structure, reinforcing its effectiveness in underwater image enhancement tasks.
@@ 5 Conclusion
In this work, we presented FUSION: Frequency-guided Underwater Spatial Image recOnstructioN, a novel dual-domain framework for underwater image enhancement. By integrating multi-scale spatial feature extraction with adaptive attention mechanisms and complementing it with frequency domain processing via FFT-based transformations, we effectively address complex degradations inherent to underwater environments.
Our experimental evaluations on the UIEB, EUVP, and SUIM-E datasets demonstrate that FUSION outperforms state-of-the-art methods in terms of reconstruction fidelity, perceptual quality, and color correction, as evidenced by significant improvements in PSNR, SSIM, LPIPS, UIQM, UISM, and BRISQUE metrics. The efficiency analysis reveals that FUSION achieves a favorable trade-off between computational cost and enhancement performance, making it well-suited for deployment on resource-constrained platforms such as autonomous underwater vehicles (AUVs).
Overall, FUSION provides an accurate, efficient, and a SOTA solution for underwater image enhancement, advancing over all the current state-of-the-art methods.
@@
Supplementary Results
In this supplementary document, we provide additional quantitative visualizations to further illustrate the performance and efficiency of our proposed FUSION framework. In addition to the primary metrics presented in the main paper, these supplementary results include detailed bar plots comparing individual quality metrics across the UIEB, EUVP, and SUIM-E datasets, as well as extended efficiency analyses. These visualizations are closely tied to the mathematical formulations described in Section 3 and further underscore the effectiveness of our dual-domain processing.
@@
To provide a granular view of the performance across different metrics, we present bar plots (Figures 11-16)for each quality measure across the UIEB, EUVP, and SUIM-E datasets. These plots allow us to compare how various methods perform in terms of perceptual quality (BRISQUE and LPIPS), reconstruction fidelity (PSNR and SSIM), reconstruction (MSE), and overall image quality (UIQM and UISM).

### v2

@@ 2.1 Underwater Image Enhancement
Traditional methods for UIE have relied on image processing techniques such as histogram equalization, white balance adjustment, and dehazing algorithms based on physical models of light propagation underwater [4, 7, 12]. While these methods can enhance contrast and correct color casts to some extent, they generally lack adaptability to varying underwater conditions and often fail to restore fine details and textures. Li et al. proposed a dehazing and color correction method using convolutional neural networks (CNNs) that leverage the statistical properties of underwater images [12]. FUnIE-GAN and Water-Net have shown promising results by learning mappings from degraded images to their enhanced counterparts using generative adversarial networks (GANs) [22, 37].
There exists minimal literature on frequency-based methods for image enhancement, with no prior application to underwater imaging. Kersting et al. used a GAN-based approach to enhance ultra-fast PSMA-PET scans via synthetic reconstruction, showing improved detection in prostate cancer staging [11]. Liu et al. applied frequency decomposition in PID2Net for underwater descattering and denoising, though not explicitly using frequency-domain learning [23]. Li et al. fused polarization cues with wavelet-based subband processing to improve defect visibility on reflective surfaces [21]. Agaian et al. proposed transform-based enhancement using orthogonal bases like Fourier and Hadamard with quantifiable performance metrics [1]. Wang proposed a parallel frequency-domain low-light framework that decouples contrast and structure restoration [32], while Wang et al. designed FourLLIE, leveraging Fourier amplitude mappings and SNR-guided fusion for efficient low-light enhancement [31].
@@ 2.3 Shortcomings Addressed
The primary limitation in current underwater image restoration and enhancement approaches is that they focus predominantly on spatial-domain processing, overlooking the frequency domain’s ability to capture global color distributions and long-range dependencies. This omission often results in residual color imbalances and artifacts, especially under severe wavelength-dependent attenuation [33]. Additionally, certain models that are able to capture these domains (like Fine-tuned GANs) require very heavy computational power, which makes them not viable for deployment and scalability scenarios [12]. Our proposed FUSION addresses these issues through a dual-domain design that efficiently processes each color channel in both spatial and frequency domains while having a quick inference time and low-memory compute. By incorporating multi-stage and multi-domain attention mechanisms with channel-wise recalibration, FUSION also preserves fine details, reduces artifacts, and balances color distributions.
@@ 3 Proposed Method: FUSION
In the spatial domain, each channel undergoes multi-scale convolution with kernel sizes $3\times 3$, $5\times 5$, and $7\times 7$ to capture features at varying receptive fields. This ensures that color-specific distortions are addressed independently, preventing the propagation of noisy features while preserving crucial wavelength-driven contextual information, as suggested in [29]. These features are refined using a Channel and Spatial Attention Module (CBAM) and residual connections to preserve information.
@@ 3.4 Inter-Channel Fusion and Channel Calibration
Figure 4: Visual comparisons on the UIEB dataset.
Figure 5: Visual comparisons on the EUVP dataset.
@@ 4 Results
Experimental Settings
Table 1: Evaluation on UIEB test set with the best-published works for UIE. First, second, and third best performances are represented in red, blue, and green colors, respectively. ↓ indicates lower is better.
@@
Table 2: Ablation hardware comparisons with respect to average performance across datasets ($\overline{\mathrm{Metric}}$ denotes the average of that metric across the 3 datasets used).
Configuration Freq. Attn Freq. Branch Freq. Fusion Chan. Calib Local Attn Global Attn Inference Time (ms) GFLOPs $\overline{\mathrm{UISM}}$ $\overline{\mathrm{LPIPS}}$ $\overline{\mathrm{BRISQUE}}$
Full Model (FUSION) 128.68 36.73 7.385 0.135 23.797
No Frequency Attention 128.53 36.71 6.395 0.207 27.643
No Frequency Branch 88.89 36.55 5.996 0.255 29.553
No Frequency Guided Fusion 90.29 36.71 6.192 0.226 28.370
No Channel Calibration 128.70 36.73 6.164 0.230 28.517
No Local Attention 75.87 36.69 6.453 0.210 27.627
No Global Attention 110.69 36.72 6.561 0.200 27.167
Spatial Only 89.01 36.55 5.908 0.250 29.320
Minimal Model 18.49 36.49 5.704 0.276 30.607
Table 3: Evaluation on EUVP dataset with the best-published works for UIE. First, second, and third best performances are represented in red, blue, and green colors, respectively. ↓ indicates lower is better.
@@
We first evaluate the performance of our proposed FUSION framework on three widely used underwater image datasets: UIEB [14], EUVP [10], and SUIM-E [28]. All images across these datasets are resized to a uniform resolution of 256×256 prior to training and evaluation. For training, we utilize the EUVP dataset, which contains 11,435 paired underwater images, while its test set consists of 515 image pairs of the same resolution. The UIEB dataset comprises 890 paired images, from which 800 are randomly selected for training, and the remaining 90 images are used for testing (following [14]). The SUIM-E dataset includes 1,635 images, with 1,525 used for training and 110 for evaluation (following [27]).
To comprehensively assess the visual quality and perceptual fidelity of enhanced images, we compare our method against a range of state-of-the-art (SOTA) underwater image enhancement (UIE) approaches using both full-reference and no-reference metrics. These include Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS), along with perceptual quality measures such as the Underwater Image Quality Measure (UIQM), Underwater Image Sharpness Measure (UISM), and Blind/Referenceless Image Spatial Quality Evaluator (BRISQUE). Tables 1 and 3 present a detailed comparison of the quantitative results on the UIEB and EUVP datasets, respectively.
4.1 Comparison with State-of-the-Art
We present a quantitative and quantitative evaluation demonstrating that FUSION consistently outperforms competing methods across all evaluated metrics, achieving state-of-the-art results. In particular, on the UIEB test set (Table 1), FUSION achieves a PSNR of 23.717 dB and an SSIM of 0.883, indicating a very high reconstruction fidelity and structural similarity. It also records the lowest LPIPS score (0.112), reflecting superior perceptual quality and detail preservation. We observe similar trends on the EUVP dataset (Table 3), where FUSION attains a PSNR of 28.671 dB and the highest SSIM value of 0.862, alongside a low LPIPS score (0.174). Figure 6 depicts a bubble chart illustrating the trade-off between average PSNR and GFLOPs for various models, further validating the balance between efficiency and effectiveness of our approach.
@@ 4.1 Comparison with State-of-the-Art
WaterNet [15] 24.8 193.7
UGAN [6] 57.17 18.3
FUnIE-GAN [10] 7.71 10.7
Ucolor [20] 157.4 443.9
SGUIE-Net [28] 18.55 123.5
DWNet [29] 0.48 18.2
Ushape [24] 65.6 66.2
LitNet [27] 0.54 17.8
@@
We evaluate the visual quality of our FUSION framework through qualitative comparisons. Figures 4 and 5 show enhancement results for the UIEB and EUVP datasets alongside outputs from state-of-the-art methods. FUSION recovers finer structural details and preserves subtle textures, restoring balanced color distributions and improving contrast to mitigate underwater distortions like color casts and low visibility. UIEB and EUVP results (Figure 4) enhance natural hues and recover important scene details better than competing methods.
@@
Table 5: Ablation performance on UIEB.
Configuration Freq. Attn Freq. Branch Freq. Fusion Chan. Calib Local Attn Global Attn UIQM UISM LPIPS BRISQUE
Full Model (FUSION) 3.414 7.429 0.112 23.19
No frequency attention 2.978 7.235 0.153 24.81
No Frequency Branch 2.903 6.606 0.231 27.25
No Frequency Guided Fusion 2.961 6.821 0.202 26.34
No Channel Calibration 2.827 6.751 0.214 26.68
No Local Attention 3.005 7.102 0.169 25.22
No Global Attention 3.000 7.268 0.148 24.37
Spatial Only 2.896 6.660 0.225 26.91
Minimal Model 2.720 6.410 0.258 28.43
Table 6: Ablation Study Results on the EUVP Dataset
Configuration Freq. Attn Freq. Branch Freq. Fusion Chan. Calib Local Attn Global Attn UIQM UISM LPIPS BRISQUE
Full Model (FUSION) 3.220 7.048 0.174 29.547
No Frequency Attention 2.839 6.118 0.227 34.21
No Frequency Branch 2.674 5.709 0.249 35.68
No Frequency Guided Fusion 2.665 5.744 0.247 35.53
No Channel Calibration 2.640 5.646 0.252 35.89
No Local Attention 2.538 6.222 0.232 34.51
No Global Attention 2.640 6.392 0.224 33.92
Spatial Only 2.373 5.557 0.261 36.43
Minimal Model 2.106 5.553 0.278 37.21
In addition to quantitative performance, we also assess the efficiency of our approach. Table 4 summarizes the model parameters and GFLOPs for our method compared to other leading UIE models at an input size of 256$\times$256. Notably, FUSION achieves superior enhancement results with a significantly lower number of parameters (0.28M) and competitive GFLOPs (36.73), justifying its potential for deployment in real-time and resource-constrained settings.
4.2 Ablation Study
Quantitative Analysis. From the ablation studies across UIEB and EUVP, it is evident that each architectural component contributes meaningfully to overall performance. Removing frequency attention, branch, or guided fusion consistently leads to notable degradation in perceptual quality (higher LPIPS, lower UIQM and UISM), affirming the critical role of frequency-aware design in FUSION. Similarly, channel calibration and attention blocks - both local and global - also drive significant gains, especially in structural sharpness and perceptual realism. Interestingly, global attention appears to be particularly vital in retaining fine-grained global coherence, while local attention improves texture fidelity. Models stripped of frequency modules or reduced to spatial-only designs suffer from reduced enhancement quality, confirming the synergy between spectral and spatial representations in underwater image enhancement.
Hardware Efficiency. Beyond accuracy, FUSION maintains competitive inference efficiency, showcasing a balanced trade-off between performance and resource footprint. The full model runs at 128.68 ms with just 36.73 GFLOPs, which is notably efficient given its multi-branch design. Ablating the frequency branch or removing attention mechanisms reduces inference time - e.g., down to 75.87 ms without local attention - but at the cost of performance. While the minimal model is fastest at 18.49 ms, it offers the weakest performance, backing the need for our architectural complexity to achieve enhancement fidelity. Overall, FUSION demonstrates that strategic architectural additions, particularly those exploiting frequency and attention cues, yield meaningful gains without sacrificing deployability in real-time or resource-limited scenarios.
@@ 5 Conclusion
We propose FUSION (Frequency-guided Underwater Spatial Image recOnstructioN), a novel dual-domain framework that combines multi-scale spatial feature extraction with FFT-based frequency processing for underwater image enhancement. Leveraging adaptive attention, FUSION effectively addresses complex degradations in underwater scenes. Extensive evaluations on UIEB, EUVP, and SUIM-E show superior performance across PSNR, SSIM, LPIPS, UIQM, UISM, and BRISQUE metrics. FUSION also offers a strong balance between quality and efficiency, making it ideal for real-time use on AUVs.
@@
Hardware and Training Details
We run all our experiments on a NVIDIA Tesla P100 GPU (Pascal architecture) with 16 GB of HBM2 memory and 3,584 CUDA cores, delivering up to 9.3 TFLOPS of single-precision performance. Since our method focuses on lightweight design and real-time feasibility, testing on a GPU with minimal compute ensures efficiency without relying on heavy hardware. We also use automatic mixed-precision (AMP) to speed up training and reduce memory usage, making the process even more efficient.
Training Settings: We train our model using the Adam optimizer with a starting learning rate of $2\times 10^{-4}$, $\beta_{1}=0.5$, and $\beta_{2}=0.999$. Training runs for up to 1,000 epochs with a batch size of 4, but we use early stopping based on LPIPS. We choose LPIPS since it closely aligns with human perception, ensuring that the model focuses on producing visually improved underwater images.
Computation-Related Info:
===== Model Performance Report =====
GPU Memory Used: 50.46 MB
Peak GPU Memory: 260.34 MB
Inference Time: 1.8147 seconds
Estimated FPS: 0.55 frames per second
Total FLOPs: 18.41 GFLOPs
Supplementary Results
In these supplementary results, we provide additional quantitative and qualitative visualizations to further illustrate the performance and efficiency of our proposed FUSION framework. In addition to the primary metrics presented in the main paper, these supplementary results include detailed ablation studies, bar plots comparing quality metrics across the UIEB, EUVP, and SUIM-E datasets, as well as extended efficiency analyses. These visualizations are closely tied to the mathematical formulations described in Section 3 and underscore the importance of our dual-domain processing.
Table 7: Evaluation on SUIM-E test set with the best-published works for UIE. First, second, and third best performances are represented in red, blue, and green colors, respectively. ↓ indicates lower is better.
Method PSNR SSIM LPIPS↓ UIQM UISM BRISQUE↓
UDCP [5] 12.074 0.513 0.270 1.648 7.537 22.788
GBdehaze [13] 14.339 0.599 0.355 2.255 7.400 20.175
IBLA [26] 18.024 0.685 0.209 1.826 7.341 20.957
ULAP [30] 19.148 0.744 0.231 2.115 7.475 21.250
CBF [2] 20.395 0.834 0.194 3.003 7.360 21.115
UGAN [6] 24.704 0.826 0.190 2.894 7.175 20.288
UGAN-P [6] 25.050 0.827 0.188 2.901 7.184 18.768
FUnIE-GAN [10] 23.590 0.825 0.189 2.918 7.121 22.560
SGUIE-Net [28] 25.987 0.857 0.153 2.637 7.090 25.927
DWNet [29] 24.850 0.861 0.133 2.707 7.381 20.757
Ushape [24] 22.647 0.783 0.213 2.873 7.061 22.876
Lit-Net [27] 25.117 0.884 0.118 2.918 7.368 19.602
FUSION (Ours) 25.989 0.850 0.118 3.183 7.679 18.655
Figure 9: Ablation Study Visual Comparisons. This figure displays the enhancement results for a representative underwater image using our model with various component ablations: Original, No Frequency Attention, No Frequency Branch, No Frequency Guided Fusion, No Channel Calibration, No Local Attention, No Global Attention, and Spatial Only. The qualitative differences underscore the contribution of each module to the final enhancement quality.
The above figure (9) illustrates the impact of ablating key components of our proposed model. We observe that removing individual modules results in visible degradations, which verifies the necessity of each part for achieving optimal performance.
Figure 10: Comparison with Traditional Image Processing Techniques. The figure compares the original underwater image with images processed by conventional methods: histogram equalization, CLAHE, white balance, gamma correction, dark channel prior, underwater dark channel prior, and red channel prior. These comparisons highlight the limitations of traditional methods relative to our approach.
Figure 10 presents a qualitative comparison between our FUSION framework and several traditional image processing techniques. Notably, while methods such as histogram equalization and dark channel priors provide some level of enhancement, they fall short of recovering natural color balance and structural details, as achieved by our method.
Figure 11: Visual comparisons on the SUIM-E dataset.
89
Figures 7 and 8 plot PSNR and SSIM values across the datasets. The PSNR chart shows FUSION consistently achieves higher reconstruction fidelity with elevated PSNR values. The SSIM chart reveals superior structural similarity compared to other approaches, even under challenging conditions. These plots highlight that FUSION enhances local details and color balance while preserving global image structure, reinforcing its effectiveness in underwater image enhancement tasks.
Table 8: Ablation performance on SUIM-E.
Config Freq. Attn Freq. Branch Freq. Fusion Chan. Calib Local Attn Global Attn UIQM UISM LPIPS BRISQUE
Full Model (FUSION) 3.183 7.679 0.118 18.655
no_frequency attention 2.626 5.832 0.242 23.91
no_frequency branch 2.806 5.674 0.285 25.73
no_frequency guided fusion 2.703 6.010 0.230 23.24
no_channel calibration 2.721 6.096 0.225 22.98
no_local attention 2.736 6.034 0.228 23.15
no_global attention 2.746 6.023 0.229 23.21
spatial only 2.645 5.506 0.265 24.62
minimal model 2.445 5.149 0.292 26.18
On the SUIM-E test set (Table 7), our approach further confirms its robustness by achieving comparable PSNR, SSIM, and LPIPS scores. Additionally, FUSION exhibits favorable performance in perceptual quality metrics, with competitive UIQM, UISM, and BRISQUE scores across all datasets.
@@
To provide a granular view of the performance across different metrics, we present bar plots (Figures 13-18)for each quality measure across the UIEB, EUVP, and SUIM-E datasets. These plots allow us to compare how various methods perform in terms of perceptual quality (BRISQUE and LPIPS), reconstruction fidelity (PSNR and SSIM), reconstruction (MSE), and overall image quality (UIQM and UISM).